If the average American driver goes 5000 mi each year on interstate highways, how much less time did the average driver spend on interstate highways each year as a result of the 1995 increase in the speed limit from 55 mi/h to 65 mi/h?
1
Expert's answer
2016-01-19T08:28:40-0500

1. Let's find how much time average American driver used to spend on interstate before 1995:

t = 5000 / 55 = 90.9 hours

2. And the same after 1995:

t = 5000 / 65 = 76.9 hours

3. And finally, the difference between time spent:

dt = 90.9 - 76.9 = 14 hours

So average American driver spends each year 14 hours less then he had spent before 1995.